# HyViLM: Enhancing Fine-Grained Recognition with a Hybrid Encoder for Vision-Language Models

### v2

@@ 4.4 Ablation studies
Table 4: Ablation study of The Optimal Interaction Structure. “Multi Layers” refers to the interaction between the lower-level features of ConvNeXt and the lower-level features of ViT, as well as the interaction between higher-level features and higher-level features. “Last Layer” means using only the final layer features of ConvNeXt to interact with ViT. “Pyramid Fusion” refers to first fusing the multi-level features of ConvNeXt, and then using the fused features as a whole to interact with ViT. “ViT Layer” indicates the number of interaction layers set in ViT.
@@
Table 5: Ablation Study of Interaction Methods. “Int. Mode” means the interaction mode between ViT and ConvNeXt in CVMF, and “CA” refers to Cross Attention. Note that the experiments here are all based on deep feature interactions.

### v3

@@ 4.4 Ablation studies
Table 3: Ablation study of the optimal interaction structure. ”Multi Layers” refers to the interaction between lower- and higher-level features of ConvNeXt and ViT. ”Last Layer” uses only the final layer features of ConvNeXt for interaction. ”Pyramid Fusion” fuses multi-level features of ConvNeXt before interacting with ViT. ”ViT Layer” indicates the number of interaction layers in ViT.
@@
Table 4: Ablation Study of Interaction Methods. Int. Mode means the interaction mode between ViT and ConvNeXt in CVMF, and CA refers to Cross Attention. Note that the experiments here are all based on deep feature interactions.
@@ 5 Conclution
Appendix A More Visualization
We selected images from DocVQA, WikiTableQuestions, and TextVQA, designed questions, and shifted the images so that the answers are at the edge of the cropped boundary. We compared our model with Minigemini-HD and LLaVA-NeXT. As shown in Figure 5, our model’s performance in answering at the cropped boundary is mostly correct, while both Minigemini-HD and LLaVA-NeXT provided incorrect answers.
Figure 5: Qualitative results. The red solid line in the image represents the cropping boundary, and the blue box indicates the location of the answer. All images are sized at $672\times 672$. Neither the red solid line nor the blue box will be input into the model.
Appendix B Training Settings
In this study, we use the ViT-L pretrained by CLIP (Radford et al., 2021) and ConvNeXt-L pretrained by LAION (Schuhmann et al., 2022) as the high-resolution visual encoder and the main architecture of Hybrid Encoder, respectively. For the language model, we utilize LLaMA3-8B-Instruct (Touvron et al., 2023). We optimize the model for 1 epoch using the AdamW optimizer with a cosine learning rate schedule and a warmup ratio of 0.03. During the pretraining stage, we use a peak learning rate of 1e-3 and a batch size of 256, training the projection and interaction layers. In the instruction fine-tuning phase, our main peak learning rate is 1e-5 with a batch size of 64. The learning rate for the visual encoder is reduced to 2e-6 to ensure stability, while the interaction layer remains at 1e-5. We employ the DeepSpeed Zero (Rajbhandari et al., 2020) 2 strategy, completing the optimization in a total of 32 hours on 32 $\times$ A800 GPUs.
Appendix C Alignment Strategy
1 (192, 192, 192) (3072, 24, 24) 2 (1024, 24, 24)
2 (384, 96, 96) (1536, 24, 24) 6 (1024, 24, 24)
3 (768, 48, 48) (3072, 24, 24) 12 (1024, 24, 24)
4 (1536, 24, 24) (1536, 24, 24) 20 (1024, 24, 24)
In the ConvNeXT-ViT Deep Fusion Module, we opted for channel concatenation. Table 5 provides a detailed illustration of the dimensional changes in the features at different stages of ConvNeXT and the corresponding interaction layers with CLIP-ViT. Here, we use an image with a resolution of $372\times 372$ as an example. According to the cropping strategy, the resolution of the images input into ConvNeXT will be $768\times 768$.